OpenStudy (youngnephew34):

A road is in the shape of a regular heptagon. What is the measure of each angle on the sign? Round to the nearest tenth? A) 128.6 B) 900 C) 64.3 D) 231.4 Can anybody help me check my work PLZZZZZZZZZ FAN and MEADAL

OpenStudy (anonymous):

what do you think the answer is

OpenStudy (youngnephew34):

I think it is A but not for sure

OpenStudy (anonymous):

sum of external angles of a regular polygon =360 one external angle of regular heptagon=360/7=51.4 internal angle=180-51.4=128.6

OpenStudy (anonymous):

so yes you are right

OpenStudy (youngnephew34):

ok